OpenStudy (anonymous):

Is the following system consistent or inconsistent? 2y = x - 7 -2x - 6y = -14 A. Consistent B. Inconsistent

OpenStudy (anonymous):

Thanks

OpenStudy (texaschic101):

if the equations have no solution, then they are inconsistent. If the equations have one or more solutions, then they are consistent. 2y = x - 7 x = 2y + 7 -2x - 6y = -14 -2(2y + 7) - 6y = -14 -4y - 14 - 6y = -14 -4y - 6y = -14 + 14 -10y = 0 y = 0 2y = x - 7 2(0) = x - 7 0 = x - 7 x = 7 check.. -2x - 6y = -14 -2(7) - 6(0) = -14 -14 - 0 = -14 -14 = -14 (correct) The solution to this system of equations is (7,0) And since it has one solution, then it is consistent

OpenStudy (texaschic101):

or you could have done this... put both equations in y = mx + b form 2y = x - 7 y = 1/2x - 7/2 slope is 1/2 and y intercept is -7/2 -2x - 6y = -14 -6y = 2x - 14 y = -1/3x + 7/3 slope is -1/3 and y intercept is 7/3 Now remember this... if they have same slopes, but different y intercepts, they are parallel lines and do not intersect. They have no solutions. if they have same slopes and same y intercepts, they are same line and have infinite solutions. If they have different slopes and different y intercepts, they will have 1 solution and be consistent